OpenStudy (anonymous):

let R^2 have the Euclidean inner product. Use the Gram-Schmidt process to transform the basis {u1,u2} into an orthonormal basis. u1= (1,-3), u2= (2,2)

OpenStudy (anonymous):

so, if we consider u1 as one basis, we want one that is orthonormal to u1

OpenStudy (anonymous):

i.e., project u2 on a perpendicular vector.

OpenStudy (anonymous):

yea. but i confuse when the answer is not that.

OpenStudy (anonymous):

you can start with u2 as your original vector too

OpenStudy (anonymous):

and find the projection of u1 on an orthogonal verctor

OpenStudy (anonymous):

sometimes, the process tells me keep v1 =u1, and count v2= ...... but in this case, they turn v1 to orthonormal vector, too

OpenStudy (anonymous):

oh ya, first count them by using Gram, second, turn to orthonomal. sorry kid.

OpenStudy (anonymous):

is it right?

OpenStudy (anonymous):

let me see, did they take \[u_1=\left[{1\over2},{-3\over2}\right]^T\]

OpenStudy (anonymous):

don't get!!

OpenStudy (anonymous):

normalized the vector!

OpenStudy (anonymous):

yea.. my bad \[ u_1=\left[{1\over\sqrt{10}},-{3\over\sqrt{10}}\right]^T \]

OpenStudy (anonymous):

yeap

OpenStudy (anonymous):

but not T, just in brake

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what do you mean by that ok? I 'm not done yet. my question is: getting v1 and count v2 base on that v1 or taking v1 , orthonormal v1 and then count v2 base on the orthonormal v1? which one?

OpenStudy (anonymous):

so, starting with v1, we find v2: \[ v_2=(2,2)^T-\frac{(2,2)\cdot(1,-3)}{(1,-3)\cdot(1,-3)}(1,-3)^T\\ v_2=(2,2)^T-\frac{-4}{10}(1,-3)^T=\left(2+{2\over5},2-{6\over 5}\right)^T\\ v_2=\left({12\over5},{4\over5}\right)^T \]

OpenStudy (anonymous):

now, normalize it

OpenStudy (anonymous):

normalize both them is the last step?

OpenStudy (anonymous):

yes, got it. thank you, sir

OpenStudy (anonymous):

yep. you cannot normalize v1 before coz you will loose the length information.

OpenStudy (anonymous):

now, i am done with it.

OpenStudy (anonymous):

you are welcome madam.